OpenStudy (mendicant_bias):

I'm trying to find the line tangent to a parametrized function and I am doing that by eliminating the parameter rather than taking the derivative of both y(t) and x(t) and dividing, but I'm doing something wrong, and I can't figure out what.

OpenStudy (mendicant_bias):

\[x(t) = 5t, \ y(t) = \sqrt{t}\]

OpenStudy (mendicant_bias):

\[\frac {x}{5} = t\] \[\sqrt{\frac {x}{5}} = \sqrt{t}\]

OpenStudy (mendicant_bias):

\[\frac {\sqrt{x}}{\sqrt{5}} = \sqrt{t} = y\]

OpenStudy (mendicant_bias):

\[\frac {dy}{dx} = \frac {1}{\sqrt{5}} \frac {d}{dx} (\sqrt{x})\]

OpenStudy (mendicant_bias):

\[\frac {dy}{dx} = \frac {1}{\sqrt{5}}* \frac{1}{2 \sqrt{x}} = \frac {1}{2 \sqrt{5} \sqrt{x}}\]

ganeshie8 (ganeshie8):

looks good ?

OpenStudy (mendicant_bias):

Yeah, it says that it should be:

OpenStudy (mendicant_bias):

Because they chose not to eliminate the parameter, which I haven't taken a look at, but the math should be the same result

OpenStudy (mendicant_bias):

lemme post the whole question verbatim just to make sure this isn't a bookkeeping issue

ganeshie8 (ganeshie8):

yup ! \(\large \frac{dy}{dx} = \frac{dy}{dt} \frac{dt}{dx} = \frac{dy}{dt} / \frac{dx}{dt} \)

OpenStudy (mendicant_bias):

Huh. Guess I'll let me prof. know about this.

ganeshie8 (ganeshie8):

about what ?

OpenStudy (mendicant_bias):

Lemme go through the parametrized version just to make sure. About this not matching up. IDK why they wouldn't be the same answer.

ganeshie8 (ganeshie8):

you wud get the same answer

OpenStudy (mendicant_bias):

Oh, derp.

ganeshie8 (ganeshie8):

we just need to replace "t" to "x" in the end :)

OpenStudy (mendicant_bias):

Yeah, lol...thanks

ganeshie8 (ganeshie8):

the second derivative is a bit trickier in parametric form

OpenStudy (mendicant_bias):

I'll try to do it myself, as long as I don't make any clumsy errors, I hopefully will be able to do it :3 I'll just leave it in cartesian form. ty.

ganeshie8 (ganeshie8):

you will be using parametric form a lot in multivariable calculus, so it is better to do it in parametric form.. gives some practicce :)

ganeshie8 (ganeshie8):

cartesian form is straightforward.. .

ganeshie8 (ganeshie8):

good luck !

OpenStudy (mendicant_bias):

\[\frac {dy^{2}}{d^{2}x} = \frac {1}{2 \sqrt{5}} \frac {d}{dx}(\sqrt {x}) = - \frac {1}{4 \sqrt{5}x^{3/2}}\] \[\frac {dy^{2}}{d^{2}x} = \frac {d}{dt} \left( \dfrac {dy}{dx} \right) / \frac {dx}{dt} = -\frac {1}{20 t^{3/2}} / \ 5 = - \frac {1}{100t^{3/2}}\]

OpenStudy (mendicant_bias):

(Thanks, btw!)